A cylindrical diving Bell whose length is 150 cm is lowered to the bottom of a tank. The water is found to write 50 cm in the bell. Find the depth of the tank. Assume the atmospheric pressure at the surface is equivalent to 100 cm height of water and the temperature as constant

Grade:11

1 Answers

Arun
3 years ago
pressure at surface =1000x150a
pressure in water=100ax(1000-50+h)
P1v1=p2v2
(950+h)100=1000x150
h=1500-950=550